# MSegRNN:Enhanced SegRNN Model with Mamba for Long-Term Time Series Forecasting

### v2

@@ 5 Experiments
Our experiments are structured as follows. In Section 5.1, we introduce the datasets used for long-term time series analysis. Subsequently, in Section 5.2, we provide a detailed description of the experimental setup and baseline models for comparison. Section 5.3 presents and analyzes the performance metrics. In Section 5.4, we conducted ablation studies to investigate the effectiveness of the model. Finally in Section 5.5, we investigated the efficiency of the model, particularly focusing on its memory consumption to demonstrate the lightweight characteristic of our model. All experiments in this section were conducted on two NVIDIA T4 GPU, The source code is available at https://github.com/UniqueoneZ/MSegRNN.git.
@@ 5.1 Datasets
The evaluation is conducted across seven real-world datasets spanning various domains, including power transformer temperature, weather, electricity load, and traffic, with channel ranging from 7 to 321 and frequencies from every 10 minutes to hourly recordings. Detailed dataset information is provided in Table 1.
@@
Dataset Electricity Weather ETTh1 ETTh2 ETTm1 ETTm2
Time-Series 321 21 7 7 7 7
Frequency 1 hour 10 mins 1 hour 1 hour 15 mins 15 mins
Time-Points 26,304 52,696 17,420 17,420 69,680 69,680
@@ 5.3 Main Result
Our method achieved top positions in 36 out of 48 metrics, while the SegRNN method achieved the top position in 7 metrics. The iTransformer, Crossformer, and PatchTST methods each achieved top position in 6, 1 and 1 metrics, respectively. This demonstrates the powerful capabilities of our model.
@@
Datasets Models RNNs Transformers MLPs CNNs
@@ 5.4 Ablation study
We further conducted an ablation study to verify the effectiveness of incorporating the Mamba structure, implicit segmentation, and residual structures into our model. Using the weather dataset as an example, we performed experiments excluding the Mamba structure and implicit segmentation structure respectivly. For clarity, in the table below, "M" denotes preprocessing using the Mamba structure, and "LR" represents the addition of linear segmentation and residual structure. The results are shown in Table 3:
@@
Average 0.335 0.355 0.341 0.357 0.425 0.411 0.340 0.372
The ablation study reveals that the model incorporating the Mamba structure, along with implicit segmentation and residual structures performs the best overall. This is followed by the model that includes only implicit segmentation and residual structures. Interestingly, merely adding the Mamba structure does not significantly enhance model performance. This may be due to a synergistic effect between the two structures. Consequently, the ablation study confirms the superiority of our proposed structure.
@@ 5.5 Model efficiency
Since our Mamba structure has not undergone hardware-level optimization, this section focuses on comparing the model’s memory usage to verify its lightweight characteristic. The lightweight characteristic is primarily due to the use of a small number of parameters in our Mamba structure experiments. Using the Weather dataset as an example, the memory usage comparison with batch size 64 is shown in Table 4.
Table 4: Memory Usage of Two Methods
Method Memory Usage (GB)
96 192 336 720
MSegRNN 0.98 1.17 1.56 2.48
SegRNN 0.91 1.14 1.53 2.41
Compare +7.7% +2.6% +2.0% +2.9%
As shown in the table, the model’s exhibits only marginal increases in memory usage compared to SegRNN, affirming its lightweight characteristics.
@@ Appendix B Code Framework
The implementation of our model framework is primarily derived from the official SegRNN code and the minimal implementation of Mamba. This includes data preprocessing and evaluation, model parameter settings and device selection, as well as result output. We sincerely thank the authors of these open-source frameworks and methods for their contributions to our experiments.
The official SegRNN code is available at:https://github.com/lss-1138/SegRNN.
The official minimal implementation of Mamba can be found at:https://github.com/johnma2006/mamba-minimal
@@ Appendix C Detail Configuration
Datasets d_state seg_len d_model learning_rate dropout batch_size
ETTh1 2 12 512 0.0003 0.1 256
ETTh2 4 24 512 0.0015 0.5 256
ETTm1 4 24 512 0.0010 0.1 256
ETTm2 4 24 512 0.0010 0.1 256
Weather 4 24 512 0.0007 0.1 64
Electricity 2 48 512 0.0012 0.01 128
@@ Appendix D Comparison with extra methods
Table 6: Perfromance of different models

### v3

@@ 4 Experiments
Our experiments are structured as follows. We first introduce the datasets used for long-term time series analysis. Subsequently, we provide a detailed description of the experimental setup and baseline models for comparison. Then we presents and analyzes the performance metrics. We further conducted ablation studies to investigate the effectiveness of the model. Finally, we investigated the efficiency of the model, particularly focusing on its memory consumption to demonstrate its lightweight characteristic. All experiments in this section were conducted on two NVIDIA T4 GPU, The source code is available at https://github.com/UniqueoneZ/MSegRNN.git.
@@ 4.1 Datasets
The evaluation is conducted across seven real-world datasets spanning various domains, including power transformer temperature, weather, electricity load, with channel ranging from 7 to 321 and frequencies from every 10 minutes to hourly recordings. Detailed dataset information is provided in Table 1.
@@
Data Electricity Weather ETTh1 ETTh2 ETTm1 ETTm2
Chan. 321 21 7 7 7 7
Freq. 1 hour 10 mins 1 hour 1 hour 15 mins 15 mins
Points 26,304 52,696 17,420 17,420 69,680 69,680
@@ 4.3 Main Result
Our method achieved top positions in 36 out of 48 metrics, while the SegRNN method and the iTransformer, each achieved top position in 7 and 6 metrics, respectively. This demonstrates the powerful capabilities of our model.
@@
Datasets RNNs Transformers MLPs CNNs
@@ 4.4 Ablation study
We further conducted an ablation study to verify the effectiveness of incorporating the Mamba structure, implicit segmentation, and residual structures into our model. We performed experiments excluding the Mamba structure and implicit segmentation structure respectivly. For clarity, in the table below, "M" denotes preprocessing using the Mamba structure, and "LR" represents the addition of linear segmentation and residual structure. The results are shown in Table 3:
@@
Average 0.334 0.355 0.341 0.357 0.425 0.411 0.340 0.372
@@ 4.5 Model efficiency
This section is dedicated to evaluating the model’s time and memory efficiency to substantiate its lightweight nature. The experiments are conducted on weather datasets because the other datasets utilize the Mamba architecture are implemented without convolutional layers and hardware optimization. The lightweight characteristic is primarily due to the use of a small number of parameters in our Mamba structure experiments. Using the Weather dataset as an example, the time and memory usage comparison with batch size 8 conducted on a single NVIDIA T4 GPU is shown in Figure 4.
Figure 4: Comparsion of SegRNN and MSegRNN with time and space efficiency.
As shown in the figure, compared to SegRNN, our model only increases memory usage and training time. However, in contrast to some transformer-based models, our method still holds significant spatiotemporal advantages.
@@ Appendix B Code Framework
The implementation of our model framework is primarily derived from the official SegRNN code and the implementation of Mamba. This includes data preprocessing and evaluation, model parameter settings and device selection, as well as result output. We sincerely thank the authors of these open-source frameworks and methods for their contributions to our experiments.
The official SegRNN source code is available at: https://github.com/lss-1138/SegRNN.
The official minimal implementation of Mamba can be found at:https://github.com/johnma2006/mamba-minimal.
The official implementation of Mamba can be found at:https://github.com/state-spaces/mamba.git
@@ Appendix C Detail Configuration
Datasets d_state seg_len use_conv learning_rate dropout
ETTh1 2 12 False 0.0003 0.1
ETTh2 4 24 False 0.0015 0.5
ETTm1 4 24 False 0.0010 0.1
ETTm2 4 24 False 0.0010 0.1
Weather 4 24 True 0.0007 0.1
Electricity 2 48 False 0.0012 0.01
@@ Appendix D Comparison with extra methods
Table 5: Performance of different models
@@
Appendix E Prediction Result
We subsequently observed the partial prediction results of the model using the weather dataset, with a look-back window of 96 and a prediction horizon of 336. The results are shown in Figure 5:
Figure 5: Prediction result of different method. The four methods are Dlinear, MSegRNN, SegRNN, and PatchTST in sequence. Intuitively, our method also has the best fitting performance, which is consistent with the fact that MSegRNN has the minimum loss among the models.
Appendix F Ablation study on using convolution layers
In the main text’s model structure, we removed the convolutional layer from the Mamba structure due to performance issues. We hypothesized that local features might have been redundantly captured. To investigate the impact of the convolutional layer on the model performence, we conducted experiments. The results are shown in Table 6.
Table 6: Ablation study on convlution layer
Datasets Models WithoutConv WithConv
MSE MAE MSE MAE
96 0.365 0.384 0.412 0.426
ETTh1 192 0.415 0.414 0.444 0.443
336 0.463 0.441 0.477 0.460
720 0.468 0.460 0.502 0.492
96 0.275 0.326 0.282 0.338
ETTh2 192 0.354 0.378 0.372 0.396
336 0.404 0.416 0.406 0.424
720 0.407 0.428 0.430 0.456
96 0.314 0.345 0.323 0.355
ETTm1 192 0.361 0.374 0.373 0.384
336 0.388 0.392 0.397 0.401
720 0.448 0.430 0.452 0.432
96 0.170 0.248 0.180 0.264
Ettm2 192 0.233 0.292 0.240 0.305
336 0.294 0.331 0.299 0.343
720 0.391 0.389 0.396 0.399
It can be observed that except for the Weather dataset in our experiments, the convolutional layer may interfere with the model’s predictions in most scenarios, thereby limiting the model’s capability.